Simplify and express the result in power notation with positive exponent:
(-4)^4 ÷(-4)^8  

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

= (-4)^5/(-4)^8
= [ a^m / a^n = a^m-n ]
= (-4)^5-8
= 1/(-4)^3

Thanks